Suppose Spain can produce 1200 pounds of olives or 1600 pounds of tomatoes each week and that Italy can produce 800 pounds of olives or 800 pounds of tomatoes each week. Which country has the absolute advantage in olives? Which country has the absolute advantage in tomatoes? If the Theory of Comparative Advantage is utilized what good(s) will each country produce and how many of that good will be made by each country? What if the introduction of new agricultural technology allows Italy to increase specialized production of each good; boosting olives by an additional 800 pounds or tomatoes by 600 pounds? Now which country would have the comparative advantage in each good and show what the calculated opportunity cost is for each country/good.
Answer #1

Spain has absolute advantage in olives because it can produce more number of olives.
Spain has absolute advantage in tomatoes because it can produce more number of tomatoes.

Opportunity cost of 1 olive for Spain = Total tomatoes given up/Total olives produced = 1600/1200 = 1.33 tomato
Opportunity cost of 1 olive for Italy = Total tomatoes given up/Total olives produced = 800/800 = 1 tomato
So, Italy has comparative advantage in olives because of lower opportunity cost.
Thus, Italy will produce olives = 800 and Spain will produce tomatoes = 1600

Spain:
Total olives after new technology = 1200 + 800 = 2000
Total tomatoes after new technology = 1600 + 600 = 2200

Opportunity cost of 1 olive for Spain = Total tomatoes given up/Total olives produced = 2200/2000 = 1.1 tomato
Opportunity cost of 1 olive for Italy = Total tomatoes given up/Total olives produced = 800/800 = 1 tomato
So, Italy has comparative advantage in olives because of lower opportunity cost.

Opportunity cost of 1 tomato for Spain = Total olives given up/Total tomatoes produced = 2000/2200 = 0.91 olive
Opportunity cost of 1 tomato for Italy = Total olives given up/Total tomatoes produced = 800/800 = 1 olive
So, Spain has comparative advantage in tomatoes because of lower opportunity cost.